OpenStudy (anonymous):

5x-(7x-4)-2=5-(3x+2)

OpenStudy (anonymous):

Distribute the negative value and the add like terms then solve for x

OpenStudy (anonymous):

5x-7x+4-2=5-3x-2 -2x+2=-3x+3 x=1

hero (hero):

5x-(7x-4)-2=5-(3x+2) 5x-7x+4-2=5-3x-2 -2x+2=3-3x x = 1